OpenStudy (anonymous):

A block initially at rest is pushed horizontally along a frictionless surface with a constant force. After moving a distance d the speed of the block is v. What is the speed of the block after being pushed a distance 2d?

OpenStudy (anonymous):

 workdone by the force on the block=change in kinetic energy of the block F*d=1/2mv^2-0 m=2Fd/v^2 therefore similarly F*2d=1/2mV1^2-0 =>2Fd=1/2*(2Fd/v^2)V1^2 =>V1=sqrt(2)*v......answer